The seventh term of geometric sequence is 32 and the last term is 512. If the first term is 1/2, find the number of term and the sum of the progression
Answer by Boreal(15235)

2^(-1) is the first term, and 2^9=512
There are 11 terms
the sum is 11(1-r^n)/(1-r)
=(1/2)(1-2048/-1
=(1/2)(2047)
=1023(1/2)